Consider two aqueous solutions of nitrous acid (HNO2). Solution A has a concentration of (HNO2] = 0.25 M and solution B has aFind the pH of a 0.310 M aqueous benzoic acid solution. For benzoic acid, Ka = 6.5 • 10-5. Express your answer using two deciFind the pH of a 0.150 M solution of a weak monoprotic acid having Ka = 2.0x10-5. Express your answer to two decimal places.Find the pH of a 0.150 M solution of a weak monoprotic acid having Ka = 1.0x10-3. Express your answer to two decimal places.Find the pH of a 0.150 M solution of a weak monoprotic acid having Ka = 0.15. Express your answer to two decimal places. O AD
Answer #1

part- A 2 option b is correct The ionization of Neak and decreases with increase in concentration. This is due to tess Ho ispH of weak mono protic aug - 0.15 kaz 2x105 HA-HA- cat) cd ct de-ka = 2410 - 1.33410 0.15 d=11.337154 = 1.154152 H+ 2 ca [ht]0.15 o.15 part -D c = 0.15 Ka=0.15 = 1.5420 ar ko 1 2 = 1 L = = 1 x=1 de 1x100 = 100%, (H+] = ca = 0.1541 -0.15 [H +] = 0.15